OpenStudy (bananas1234):

Simplify: -27a^2b^2/9ab

OpenStudy (bananas1234):

@robtobey

OpenStudy (bananas1234):

@Nnesha

Nnesha (nnesha):

exponent rules ?

Nnesha (nnesha):

\[\huge\rm \frac{ -27a^2b^2 }{ 9ab }\] \[\large\rm \frac{ x^m }{ x^n } =x^{m-n}\]

Nnesha (nnesha):

u add when you multiply *same* bases

OpenStudy (bananas1234):

i can't multiply the same bases here can i?

Nnesha (nnesha):

nope bec it's a fraction

OpenStudy (bananas1234):

ok

Nnesha (nnesha):

 move exponent from denominator to de top when you do that signs chane